# PuzzleVQA: Diagnosing Multimodal Reasoning Challenges Through Abstract Patterns

### v1

@@ 8 Conclusion
9 Limitations
In this work, we mainly focus on the zero-shot setting to investigate how large multimodal models face reasoning challenges in novel situations. However, previous works have shown that prompting with demonstrations Brown et al. (2020) may improve the models ability to adapt to new tasks. Hence, we also include experiments in the one-shot in Appendix A.4, which showed inconsistent benefits, and we aim to explore this area in the future works.
@@ A.3 Prompt Examples
A.4 One-Shot Setting
In-context learning Brown et al. (2020) is a popular prompting method that can help models generalize better to tasks when they are provided with one or more task demonstrations. Hence, to investigate whether the large multimodal models are capable of learning to reason better over abstract patterns, we report the one-shot evaluation results in Table 5. Concretely, given a testing puzzle, we randomly sample a demonstration puzzle from the same concept category. To ensure that the models can learn and generalize to novel puzzles, we ensure that the testing puzzle and demonstration puzzle do not demonstrate the same abstract pattern. The model is then provided with the multimodal input of the demonstration puzzle, as well as the ground truth explanation consisting of the visual perception, inductive reasoning, and deductive reasoning.
While we observe generally better performance for Gemini Pro and GPT-4V, we note that LLaVA-13B did not improve. Furthermore, the performance does not improve consistently across all concept categories. For instance, we find that GPT-4V improved for colors and shapes, but not numbers or size. Hence, we plan to investigate how to improve the in-context reasoning ability of multimodal models in future work.
@@ A.4 One-Shot Setting
Random Baseline 25.0 25.0 33.3 25.0 27.1
LLaVA-13B 26.5 27.0 32.5 23.0 27.2
Gemini Pro 37.0 38.0 34.0 42.5 37.9
GPT-4V 65.5 54.0 30.5 47.5 49.4
Table 5: Additional results for one-shot performance on single-concept abstract patterns.

### v3

@@ 7 Conclusion
Acknowledgement
This work was substantially supported by DAMO Academy through DAMO Academy Research Intern Program. This work was partially supported by AI Singapore Governance grant ID: AISG3-GV-2023-010, and AcRF MoE Tier-2 grant (Project no. T2MOE2008, Grantor reference no. MOE-T2EP20220-0017) titled: “CSK NLP: Leveraging Commonsense Knowledge for NLP”. Chia Yew Ken would like to thank Tan Hui Min Grace as a source of inspiration for fun and unique puzzles.
Limitations
In this work, we mainly focus on the zero-shot setting to investigate how large multimodal models face reasoning challenges in novel situations. However, previous works have shown that prompting with demonstrations Brown et al. (2020) may improve the models ability to adapt to new tasks. Hence, we also include experiments in the few-shot setting in Section 5.3, which showed inconsistent benefits, and we aim to explore this area in the future works.
@@ A.3 Prompt Examples
A.4 Code Implementation Example
import math
import random
from typing import List, Tuple, Dict
from PIL import Image, ImageDraw, ImageFont
from pydantic import BaseModel
class ColorHexagonPattern(BaseModel):
colors: Dict[str, str] = dict(
blue="#6fa8dc",
green="#93c47d",
yellow="#ffd966",
red="#e06666",
purple="#8e7cc3",
orange="#f6b26b",
image_size: int = 512
scale_factor: int = 4
path_font: str = "fonts/OpenSans-Medium.ttf"
@staticmethod
def get_centroid(points: List[Tuple[float, float]]) -> Tuple[float, float]:
x = sum(p[0] for p in points) / len(points)
y = sum(p[1] for p in points) / len(points)
return x, y
def sample_colors(self) -> Tuple[List[str], List[str]]:
while True:
names = random.sample(list(self.colors), k=3)
if "orange" in names and "yellow" in names:
continue # Hard to distinguish
names = names + names
colors = [self.colors[n] for n in names]
return names, colors
def make_sample(self):
# Set the size of the image
size = self.image_size * self.scale_factor
image = Image.new("RGB", size=(size, size), color="white")
draw = ImageDraw.Draw(image)
center = size // 2
# Hexagon properties
length = size // 3 # Length of a side of the hexagon and triangles
triangle_height = math.sqrt(3) / 2 * length
# The vertices of the hexagon
hexagon = [
(center + length / 2, center - triangle_height),
(center - length / 2, center - triangle_height),
(center - length, center),
(center - length / 2, center + triangle_height),
(center + length / 2, center + triangle_height),
(center + length, center),
# Colors for the triangles
names, colors = self.sample_colors()
i_answer = random.randint(0, len(colors) - 1)
answer = names[i_answer]
colors[i_answer] = "#eeeeee" # Grey
# Draw the hexagon made of six triangles
for i in range(6):
# Coordinates of the triangle vertices
triangle = [hexagon[i], hexagon[(i + 1) % 6], (center, center)]
# Draw the triangle
draw.polygon(triangle, fill=colors[i])
# Draw the outline with custom width
points = [hexagon[i], hexagon[(i + 1) % 6], (center, center), hexagon[i]]
draw.line(points, fill="black", width=self.scale_factor * 4)
# Draw "?" on the missing answer part
if i == i_answer:
draw.text(self.get_centroid(triangle),
text="?",
font=ImageFont.truetype(self.path_font, size=size // 10),
anchor="mm",
fill="black",
names[i_answer] = "?"
instances = sorted(set(n for n in names if n not in [answer, "?"]))
image = image.resize((self.image_size, self.image_size), Image.LANCZOS)
return (
dict(
question="What is the missing color of the part denoted with a question mark?",
answer=answer,
options=sample_options(answer, options=list(self.colors), k=4),
caption=f"There is a hexagon split into six parts with the colors {names} in an anti-clockwise order.",
explanation=f"We observe that a {instances[0]} part is opposite another {instances[0]} part, and a {instances[1]} part is opposite another {instances[1]} part. Thus, the pattern is that the colors in opposite parts are the same.",
deduction=f"Based on the pattern that spatially opposite parts have the same color, the missing color of the part which is opposite a {answer} part should be {answer}.",
image,
A.5 Dataset Size Analysis
Regarding the dataset size and diversity, we set the number of generated instances for each puzzle to 100, to reduce experimental variance and maintain a reasonable evaluation cost. Hence, the current dataset size is 2000 samples (20 templates with 100 instances each). As there are two templates per puzzle category, this means there are 200 test samples for each puzzle category. To observe the impact of the number of test samples, we evaluate the models on three different data settings as shown in Table 4: 50, 100, and 200 test samples in per puzzle, which correspond to 1000, 2000, and 4000 total samples respectively. In general, we observe some variations in the average score for single-concept puzzles, but it does not significantly affect the comparison of performance between different models. Hence, we believe that the chosen dataset size is large enough to be relatively robust to experimental variance. To investigate the multimodal reasoning capabilities across diverse abstract scenarios, we construct the puzzles based on four fundamental concepts: numbers, colors, shapes, and size. Furthermore, to evaluate how well the models can relate to multiple concepts, we design both single-concept and dual-concept puzzles, and the taxonomy of diverse puzzles is shown in Figure 3.
1000 2000 4000
Gemini Pro Avg. Score 29.2 34.5 32.2
GPT-4V Avg. Score 48.8 46.4 48.4
Table 4: Analysis of model performance with respect to number of testing data samples.
A.6 Comparison of CoT and Direct Prompting
To investigate the effect of direct prompting without chain of thought, we evaluated the models on our main setting as shown in Table 1 and 2. The results are shown in Table 5 for single-concept puzzles and indicate that direct prompting is less effective for Gemini Pro and GPT-4V models, compared to CoT prompting in Table 1. This may be due to differences in training data and alignment methods between the models.
@@ A.6 Comparison of CoT and Direct Prompting
LLaVA-13B 25.5 29.0 37.0 39.0 32.6 (+5.1)
Gemini Pro 16.0 28.5 35.0 49.0 32.1 (-2.4)
GPT-4V 72.5 26.5 30.0 45.5 43.6 (-2.8)
Table 5: Results of direct prompting and change in average performance compared to CoT prompting.
Figure 8: Case study on two sample predictions from GPT-4V. The example on the left shows visual perception failures and the example on the right shows the faulty inductive reasoning of the model which proposed a spurious pattern in the image.
Appendix B Qualitative Analysis